OpenStudy (anonymous):

solve the equation and check the solution -8(t-6)-(t+9)=4

Directrix (directrix):

@luke642 Have you worked on this problem yet? You'll need to use the distributive property. -8(t-6)-(t+9)=4 -8*(t-6) -1* (t+9) = 4 -8 (t) - 8 (-6) - 1(t) - 1(9) = 4 What comes next?

OpenStudy (anonymous):

the solution is t=? (type an integer or a simplified fraction)

Directrix (directrix):

@luke642 You need to simplify -8 (t) - 8 (-6) - 1(t) - 1(9) and then set the result equal to 4. Then, you'll solve the equation. The solution will be a Real number, most likely a fraction which you may want to convert to a decimal. I don't know what your instructions are with regard to the format of the answer. Let's get the answer first. :)

OpenStudy (anonymous):

5.8888889

Directrix (directrix):

@luke642 --> How did you get that?

OpenStudy (anonymous):

Simplifying -8(t + -6) + -1(t + -9) = 4 Reorder the terms: -8(-6 + t) + -1(t + -9) = 4 (-6 * -8 + t * -8) + -1(t + -9) = 4 (48 + -8t) + -1(t + -9) = 4 Reorder the terms: 48 + -8t + -1(-9 + t) = 4 48 + -8t + (-9 * -1 + t * -1) = 4 48 + -8t + (9 + -1t) = 4 Reorder the terms: 48 + 9 + -8t + -1t = 4 Combine like terms: 48 + 9 = 57 57 + -8t + -1t = 4 Combine like terms: -8t + -1t = -9t 57 + -9t = 4 Solving 57 + -9t = 4 Solving for variable 't'. Move all terms containing t to the left, all other terms to the right. Add '-57' to each side of the equation. 57 + -57 + -9t = 4 + -57 Combine like terms: 57 + -57 = 0 0 + -9t = 4 + -57 -9t = 4 + -57 Combine like terms: 4 + -57 = -53 -9t = -53 Divide each side by '-9'. t = 5.888888889 Simplifying t = 5.888888889

OpenStudy (anonymous):

it looks wrong

Directrix (directrix):

@Luke642 --> You appear to have skipped several steps. -8 (t) - 8 (-6) - 1(t) - 1(9) = 4 -8t + 48 - t - 9 = 4 -9t + 39 = 4 Please check my work and then solve for t showing the steps. Thanks.

OpenStudy (anonymous):



OpenStudy (anonymous):

is that correct directrix? -35/9 thanks ello_bello,for breaking it down, it just seems both of you did the step the same but different order

Directrix (directrix):

@ello_bello @luke642 x = 35/9 which approximates to 3.888888... Luke, your work is detailed and thorough. I am looking for the error. Hold on. While I am checking for the error, would the two of you take the original equation and show that the answer checks by substituting 35/9 for t in: -8(t-6)-(t+9)=4

Directrix (directrix):

@luke642 The original equation is: -8(t-6)-(t+9)=4. Note that the 9 is positive-->(t+9) In your work, the 9 is negative within the parentheses -->Simplifying -8(t + -6) + -1(t + -9) = 4 After the Distributive Property is applied that -9 will be -1(-9) = 9. It should be -9. It appears that you distributed the -1 to th -9 twice.

OpenStudy (anonymous):

thanks directrix

Directrix (directrix):

Glad to help. Your work was thorough. The 9 was accidentally changed to -9 when you wrote the problem. No worries. Have you checked this problem yet?

OpenStudy (anonymous):

im taking a test so i cant check to see if its correct till im done with the whole test, but thanks

OpenStudy (anonymous):

its correct

OpenStudy (anonymous):

solve the linear inequality 5- 9/5x \[\ge6\]

OpenStudy (anonymous):

help please